Find the vertices and asymptotes of the hyperbola.
(x^2/49)+(y^2/9)=1
a.vertices:(0,+-7) asymptote:y=+-3/7x
b.vertices:(+-7,0) asymptote:y=+-7/3x
c.vertices:(+-7,0) asymptote:y=+-3/7x
d.vertices:(+-7,3) asymptote:y=+-7/3x
e.vertices:(0,+-7) asymptote:y=+-7/3x

Answer by lwsshak3(11628)

Find the vertices and asymptotes of the hyperbola.
(x^2/49)+(y^2/9)=1
a.vertices:(0,+-7) asymptote:y=+-3/7x
b.vertices:(+-7,0) asymptote:y=+-7/3x
c.vertices:(+-7,0) asymptote:y=+-3/7x
d.vertices:(+-7,3) asymptote:y=+-7/3x
e.vertices:(0,+-7) asymptote:y=+-7/3x
**
(x^2/49)+(y^2/9)=1
This is not a hyperbola: Standard form: (x-h)^2/a^2-(y-k)^2/b^2=1
What you have here is an ellipse with horizontal major axis. Ellipses do not have asymptotes.